OpenStudy (anonymous):

A car is traveling 118km/h. What is its speed in mm/s? Please explain step by step.

OpenStudy (zpupster):

this needs to be posted in math section

OpenStudy (zpupster):

but \[(\frac{ 118km }{ hr })(\frac{hr }{ 60\min })(\frac{\min }{ 60\sec })(\frac{1000m }{ km })(\frac {1000mm }{ m })\]

OpenStudy (zpupster):

units cancel and give you mm/s